OpenStudy (anonymous):

what is (9/3x)-(4/2x*) *=2

OpenStudy (anonymous):

but put it as a fraction like 9 over 3x - 4 over 2x*

OpenStudy (anonymous):

9/3x=3/x and 4/2x=2/x this means (3/x)-(2/x)=2